OpenStudy (anonymous):

John flips a fair, two-sided coin 7 times. a. What is the probability that he flips exactly 3 heads? b. What is the probability that he doesn’t get at least two heads? c. What is the probability that he doesn’t get more heads than tails? d. What is the probability that the fourth toss isn’t heads?

OpenStudy (anonymous):

if someone can just tell me how to solve each question i can do thw work myself

OpenStudy (anonymous):

*the

OpenStudy (anonymous):

@Zarkon @tcarroll010 @Hero @experimentX @RadEn can any of you help please?

OpenStudy (raden):

a). 7!/(3!4!) * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 = 35(1/2)^7

OpenStudy (anonymous):

is that first part a permutation or combination? before you multiple all the halves?

OpenStudy (raden):

it is a permutation that has with the same elements.. the events 3H's it means the rest is 4T's right ? so, use the formula : n!/(q1!q2!) with n = number elements, q1, q2 are the same elements

OpenStudy (anonymous):

ok that makes sense :)

OpenStudy (anonymous):

how do i solve b,c, and d?

OpenStudy (experimentx):

B) find the probability that he gets 6 tails + probability of 7 tails C) find the probability that (4 tails + 5 tails + 6 tails + 7 tails) D) probability that fourth coin isn't head = this seems to be single coin case

OpenStudy (experimentx):

B) find the probability that he gets 6 heads + probability of 7 heads (although results are same)

OpenStudy (raden):

B) can u find number of events (1H,6T's) + (7T's) ?

OpenStudy (anonymous):

for B wouldnt that mean it would be 6/7 + 7/7? which wouldnt make sense since you would get 1 and 6/7

OpenStudy (anonymous):

i dont understand :/

OpenStudy (raden):

case I ; many possible events for (HTTTTTT), are 7!/(6!) = 7 ways right ? case II : the event for (TTTTTTT), only 1 way right ?

OpenStudy (anonymous):

that still doesnt make sense....

OpenStudy (anonymous):

yeah i get what u mean above but what does that have to do with the answer of the probability that he doesn’t get at least two heads

OpenStudy (raden):

ok, ur question (B) ask "doesn’t get at least two heads" it means the events are : the event (1H,7T's) and the event (7T's)

OpenStudy (anonymous):

you mean (1H, 6T)?

OpenStudy (raden):

opss.. yes, sorry typo :)

OpenStudy (anonymous):

that still wont make sense with what u gave me i still have no clue how to solve it

OpenStudy (anonymous):

is there some type of formula i can use to make this way easier?

OpenStudy (anonymous):

the probability is 6/7

OpenStudy (anonymous):

of getting tails

OpenStudy (raden):

looks depending situations and conditions too :) probability = (7+1)*(1/2)^7 = 8(1/2)^7

OpenStudy (anonymous):

where did u get the 7 plus 1?

OpenStudy (raden):

sorry, if my explaination didnt make sense. maybe @Zarkon can helps to make sense :)

OpenStudy (raden):

i have done it above @LivForMusic

OpenStudy (anonymous):

oh ok i get it :)

OpenStudy (anonymous):

would C be (4/7)*(1/2)^7 ?

OpenStudy (raden):

exactly, the last is (1/2)^7 but how come u got 4/7 ? actually, it just number of events.

OpenStudy (anonymous):

would it be (4 + 5 + 6 + 7) ?

OpenStudy (anonymous):

(4 + 5 + 6 + 7)*(1/2)^7

OpenStudy (raden):

the events of "doesn’t get more heads than tails" it means the events are : (1H,6T's) = 7!/(1!6!) = 7 ways (2H,5T's) = 7!/(2!5!) = 21 ways (3H,4T's) = 7!/(3!4!) = 35 ways so, total = (7+21+35) = 63 ways thefore, the probability = 63(1/2)^7

OpenStudy (anonymous):

oh ok i keep fogetting that :/ lol. ok now D is just super confusing... i have no clue how to even start it.

OpenStudy (anonymous):

*forgetting

OpenStudy (anonymous):

i know D would be something like (???T???)

OpenStudy (raden):

yes, i think so like u @LivForMusic :)

OpenStudy (anonymous):

yeah i just dont know how i would write that probability

OpenStudy (anonymous):

would is just be 1/2?

OpenStudy (anonymous):

*it

OpenStudy (raden):

do u agree this : for the first roll, there are 2 ways (H or T) the 2nd, there are 2 ways (H or T) the 3rd, there are 2 ways (H or T) the 4th, miss it ... (have knew) the 5th, there are 2 ways (H or T) the 6th, there are 2 ways (H or T) the 7th, there are 2 ways (H or T) so, the total events = 2*2*2 * 2*2*2 = 64 ways in other words, the number all subset is 2^6 = 64 ways therefore, the probability = 64(1/2)^7

OpenStudy (anonymous):

that makes sense :) wow i wouldve never guessed to do it that way.

OpenStudy (anonymous):

i have one more question but ill post it as a new one could you help me?